# Perverse microsheaves

### v2

@@ 7.3. Perverse microsheaves on complex contact and symplectic manifolds
Being a $t$-structure is an objectwise condition. So it suffices to show, for every fixed singular Lagrangian $\Lambda$, that objects supported in $\Lambda$ satisfy the corresponding conditions; it suffices therefore to show $(\mathbb{P}\mu sh_{\Lambda,o})^{\leq 0}$, $(\mathbb{P}\mu sh_{\Lambda,o})^{\geq 0}$ define a $t$-structure on $\mathbb{P}\mu sh_{\Lambda,o}$. We may check locally. Thus we may assume $\Lambda$ is an arbitrarily small and contractible neighborhood of some point $p\in\Lambda$.
By the Kashiwara–Kawai’s general position theorem (see [13, Sec. 1.6]), we may choose a complex manifold $M$ and an embedding $\iota:\Lambda\to\mathbb{P}T^{*}M$ such that $\Lambda\to M$ is finite, hence an embedding at generic points of $\Lambda$. Let $\eta$ be the Maslov datum used on $V$; recall that the grading part is canonical. Writing $\phi$ for the fiber Maslov datum on $T^{*}M$, we use the canonical identification of the ‘grading’ parts of $\eta$ and $\iota^{*}\phi$, and any arbitrary homotopy to identify the orientation data parts (which certainly exists since $\Lambda$ is contractible). Thus we have an identification $\iota^{*}\mu sh_{\iota({\Lambda}),o}\cong\mu sh_{\Lambda,o}$. Per 5.10 and Theorem 7.10, this identification intertwines the microstalk functors used to characterize the perverse $t$-structure on the left hand side with the functors we used in Definition 5.11, which therefore defines a $t$-structure as well. ∎
We immediately deduce Theorem 1.4 from the introduction.
Proof of Theorem 1.4.
As noted in Remark 6.4, $\mathbb{C}^{*}$ acts on $W\times\mathbb{C}$ by contactomorphism fixing $W\times\{0\}$ set-wise. Hence the inclusion $({\gamma_{\mathbb{C}}})_{*}\mathbb{P}\mu sh_{W\times\{0\}}\hookrightarrow({\gamma_{\mathbb{C}}})_{*}\pi_{*}\mu sh_{\widetilde{W\times\mathbb{C}}}$ is stable under this action. Taking $\mathbb{C}^{*}$-invariants, we get a fully faithful inclusion $\mu_{\mathbb{C}}sh_{W}=(({\gamma_{\mathbb{C}}})_{*}\mathbb{P}\mu sh_{W\times\{0\}})^{\mathbb{C}^{*}}\hookrightarrow(({\gamma_{\mathbb{C}}})_{*}\pi_{*}\mu sh_{\widetilde{W\times\mathbb{C}}})^{\mathbb{C}^{*}}=({\gamma_{\mathbb{C}}})_{*}\mu sh_{W}$. The $\mathbb{C}^{*}$-action manifestly preserves the subspaces $\mu sh_{W,\mathbb{C}-c}^{\leq 0}$ and $\mu sh_{W,\mathbb{C}-c}^{\geq 0}$, so the $t$-structure passes to the invariants (taking hearts is a pullback ($\mathcal{F}^{\heartsuit}:=\mathcal{F}^{\leq 0}\cap\mathcal{F}^{\geq 0}$) hence a limit, and hence commutes with taking $G$-invariants which is also a limit). ∎
As a consequence of Theorem 7.16 and Lemma 7.6, we conclude:
Corollary 7.17.
$(\mathbb{P}\mu sh_{V,\mathbb{C}-c,o})^{\heartsuit}$ is a stack of abelian categories; in particular, the global sections $(\mathbb{P}\mu sh_{V,\mathbb{C}-c,o})^{\heartsuit}(V)$ is an abelian category.
Remark 7.18.
A $t$-structure is said to be nondegenerate if $\bigcap C^{\leq 0}=0=\bigcap C^{\geq 0}$. By co-isotropicity of microsupport, the vanishing of all microstalks implies the vanishing of an object. We conclude that if the $t$-structure on $C$ is non-degenerate, then the $t$-structure on $\mathbb{P}\mu sh_{V,\mathbb{C}-c,o}$ is also non-degenerate.
Remark 7.19.
Choose a stable (not necessarily presentable) subcategory $D\subset C$ to which the $t$-structure restricts. Require $(C,C^{\leq 0},D)$-orientation data. Then it is evident from the definitions that the truncation functors preserve, hence define a $t$-structure on, the subcategory of objects with all microstalks in $D$, characterized by the same formulas save only with e.g. $C^{\leq 0}$ replaced by $C^{\leq 0}\cap D$. For instance, we can take various bounded categories e.g. $D=C^{+}=\bigcup C^{\geq n}$, or $D=C^{-}=\bigcap C^{\leq n}$, or $D=C^{b}=C^{+}\cap C^{-}$, or ask the microstalks to be compact objects $D=C^{c}$.
Remark 7.20.
Fix $D\subset C$ as above, and assume $D^{\heartsuit}$ is Artinian. (E.g., $C=k-mod$ for a field $k$ and $D=C^{c}$.) Then the full subcategory of $(\mathbb{P}\mu sh_{V,o})^{D,\heartsuit}$ on objects with finitely stratified (rather than just locally finite) support is also Artinian. Indeed, any descending chain must have eventually stabilizing microsupports; we may restrict attention to one each in the finitely many connected components of the smooth locus of the support, hence by some point, all have stabilized.
Remark 7.21.
Let $V$ be a complex contact manifold and let $L\subset V$ be a complex Legendrian. Then $\mathbb{P}\mu sh_{V,L}(-)$ is a sheaf of stable categories while $\mathbb{P}erv_{V,L}(-)$ is a sheaf of abelian categories. Be warned however that $D(\mathbb{P}erv_{V,L}(-))$ is only a presheaf of stable categories. In particular, the natural map $D(\mathbb{P}erv_{V,L}(-))\to\mathbb{P}\mu sh_{V,L}(-)$ may restrict to an equivalence on stalks without being an equivalence on global sections. A very special case: let $V=T^{*}S^{2}\times\mathbb{\mathbb{C}}$ and $L=0_{S^{2}}\times\{0\}$. Then $\mathbb{P}\mu sh_{V,L}(L)=loc(S^{2})\otimes loc(\mathbb{C}^{*})$ while $\mathbb{P}erv_{V,L}(L)=vect_{\mathbb{C}}\otimes loc(\mathbb{C}^{*})$, due to $S^{2}$ being simply connected. Similarly, $\mu_{\mathbb{C}}sh_{T^{*}\mathbb{S}^{2},0_{S^{2}}}(0_{S^{2}})=loc(S^{2})$ while $\mu_{\mathbb{C}}sh_{T^{*}\mathbb{S}^{2},0_{S^{2}}}(0_{S^{2}})^{\heartsuit}=vect_{\mathbb{C}}$.

### v3

@@ Appendix A Existence of orientation data for real symplectic manifolds
Follows from Lemma A.4 by arguing as in Lemma 3.13. ∎
Lemma A.6.
Fix a stable quaternionic bundle $X\to BSp$. Then, any grading/orientation datum obtained from applying Proposition A.3 to the canonical grading of Definition 3.8 is canonically identified with the canonical grading/orientation datum of Definition 3.8.
Proof.
Any map $BSp\to\tau_{\leq 3}B(U/O)$ canonically factors through $\tau_{\leq 3}BSp=0$; in particular, the space of such maps is contractible. We defined the canonical grading/orientation data by taking the corresponding null-homotopy of the composition $X\to BSp\to BU\to\tau_{\leq 3}B(U/O)$.
Meanwhile the grading/orientation data from Proposition A.3 are induced by choices of null-homotopy of the map $\alpha:B\sqrt{SU}\rightarrow B^{3}(\mathbb{Z}/2)$. Any such null-homotopy induces a null-homotopy out of $BSp$ by the pre-composition
$BSp\rightarrow B\sqrt{SU}\rightarrow B^{3}(\mathbb{Z}/2)$
as explained in Lemma 3.2. ∎
Appendix B $t$-structures on Fukaya categories
Here we translate our main results across the sheaf/Fukaya correspondence of [11], in order to construct $t$-structures on Fukaya categories of certain complex exact symplectic manifolds with contracting weight 11 $\mathbb{C}^{*}$ action, such as conic symplectic resolutions and moduli of Higgs bundles.
Recall that a Liouville manifold is a (real) exact symplectic manifold $(W,\lambda)$ which is modeled at infinity on the symplectization of a contact manifold. The negative flow of the Liouville vector field $Z$ (defined by $\lambda=d\lambda(Z,\cdot)$) retracts $W$ onto a compact subset $\mathfrak{c}_{W}$ called the core. Fix if desired some larger closed conic subset $\Lambda\supset\mathfrak{c}_{W}$. We say an exact Lagrangian $L\subset W$ is admissible if it is closed and, outside a compact set, it is conic and disjoint from $\Lambda$. For example, when $Z$ is Smale and gradientlike for a Morse function, the ascending trajectories from maximal index critical points (“cocores”) are admissible. Any admissible Lagrangian disk which meets $\Lambda$ transversely at a single smooth point is termed a generalized cocore; we say a set $\{\Delta_{\alpha}\}$ of generalized cocores is complete if it meets every connected component of the smooth locus of $\Lambda$. If $Z$ is gradient-like for a Morse-Bott function, then $\Lambda$ is known to admit a complete set of generalized cocores.
Fix grading and orientation data. We recall that one source of such data is a polarization of the stable symplectic normal (or equivalently tangent) bundle as explained in Section 3 below; see also [11, Sec. 5.3]. Then one can define a partially wrapped Fukaya category $Fuk(W,\partial_{\infty}\Lambda)$ [12]. Objects are provided admissible Lagrangians equipped structures corresponding to the grading and orientation data. The object associated to a generalized cocore is unique up to grading shift. The completion of $Fuk(W,\partial_{\infty}\Lambda)$ with respect to exact triangles is generated by any any complete collection of generalized cocores [5, 12]. We further complete with respect to idempotents, and still denote the resulting category $Fuk(W,\partial_{\infty}\Lambda)$.
Definition B.1.
Fix a collection $\{\Delta_{\alpha}\}$ of generalized cocores equipped with grading data. We define:
$Fuk(W,\partial\Lambda)^{\geq 0}=\{L\,|\,\operatorname{Hom}(L,\Delta_{\alpha})\mbox{ is concentrated in degrees $\geq 0$}\}$
$Fuk(W,\partial\Lambda)^{\leq 0}=\{L\,|\,\operatorname{Hom}(L,\Delta_{\alpha})\mbox{ is concentrated in degrees $\leq 0$}\}$
It is natural to ask when Definition B.1 determines a $t$-structure.
Suppose now that $W$ is a complex manifold and that $d\lambda=\operatorname{Re}\omega_{\mathbb{C}}$, for some complex symplectic structure $\omega_{\mathbb{C}}$ on $W$. As we have remarked above, and will explain in detail in Section 3 below, such a $W$ carries canonical grading data, which agrees with the grading induced by any stable complex Lagrangian polarization of the stable complex symplectic normal bundle (viewed as a real polarization of the real stable symplectic normal bundle), and any complex Lagrangian carries a canonical secondary grading. More generally, consider a real Lagrangian (or union of Lagrangians) $L\subset W$. Then $TL$ determines a section of $LGr(TW)|_{L}$. By a $g$-complex structure on $L$, we mean a simply connected neighborhood of $LGr_{\mathbb{C}}(TW)|_{L}\subset LGr(TW)|_{L}$ containing $TL$. An $g$-complex Lagrangian has a canonical secondary grading. Evidently any Lagrangian disk transverse to a complex Lagrangian admits a canonical $g$-complex structure.
We recall the main result of [11]. Assume $W$ is real analytic and Liouville, and $\Lambda\supset\mathfrak{c}_{W}$ is subanalytic, Lagrangian at smooth points, and admits a complete collection of generalized cocores $\{\Delta_{\alpha}\}$. Fix grading and orientation data coming from a stable polarization. [^6] The dependence on polarizations here and henceforth could be removed by a version of [35, Sec. 11] for Fukaya categories. Then $Fuk(W,\partial_{\infty}\Lambda)\cong\mu sh_{\Lambda}(\Lambda)^{c,op}$ carrying $\Delta_{\alpha}$ to co-representatives of microstalk functors (here $(-)^{c}$ means that we take compact objects). Chasing definitions reveals that under [11], the normalized microstalks of Corollary 1.3 are carried to the canonically graded $\Delta_{\alpha}$ (up to some universal shift). Therefore:
Corollary B.2.
Suppose $(W,\lambda)$ is a complex exact symplectic manifold with weight 11 $\mathbb{C}^{*}$-action and $\Lambda\subset W$ a conic complex (singular) Lagrangian. Assume $(W,\operatorname{re}(\lambda))$ is Liouville and $\Lambda$ admits a complete collection of generalized cocores $\{\Delta_{\alpha}\}$.
Fix grading and orientation data induced by a stable complex Lagrangian polarization of $(W,d\lambda)$. Then the equivalence of [11] carries the $t$-structure of Theorem 1.6 to a shift of Definition B.1, which therefore provides a $t$-structure.
Also by Theorem 1.6 translated through [11], if $L,M\subset\mathfrak{c}_{X}$ are spin compact (necessarily conic) smooth Lagrangians, their Floer cohomology matches the cohomology of a (shifted) perverse sheaf supported on $L\cap M$.
The hypotheses of Corollary B.2 are obviously satisfied for $W$ a cotangent bundle of a complex manifold. More generally, there are many examples of holomorphic symplectic manifolds with a weight 11 $\mathbb{C}^{*}$ action scaling the symplectic form – (coloop-free) quiver varieties, moduli of Higgs bundles, etc. – which satisfy all the hypotheses (the Liouville flow is gradientlike for the moment map for $S^{1}\subset\mathbb{C}^{*}$, which is Morse); cf. [44].
The Fukaya category has the advantage that non-conic Lagrangians directly define objects, to which we thus have more direct geometric access. In particular, we can now construct objects in the heart. The point is that for index reasons, if $L\cup M$ is $g$-complex, then the Floer homology between $L$ and $M$ must be concentrated in degree zero. Similar considerations, for cotangent bundles, appear in [16]. Thus:
Corollary B.3.
Retain the hypotheses of Corollary B.2. Let $L\subset W$ be an exact Lagrangian. Assume $L$ is compact, or more generally, that $\partial_{\infty}L$ wraps into $\partial_{\infty}\Lambda$ without passing through any $\partial_{\infty}\Delta_{\alpha}$. Suppose $L\cup\Delta_{\alpha}$ is $g$-complex for each $\alpha$. Then $L\in Fuk(W,\partial_{\infty}\Lambda)^{\heartsuit}=\mu sh_{\Lambda}(\Lambda)^{\heartsuit}$.
The wrapping hypotheses appears to ensure that Hom in the wrapped Fukaya category is in fact computed without any wrapping, i.e. just by the Floer homology. One can imagine applying the corollary by taking an exact holomorphic Lagrangian is asymptotic to $\partial_{\infty}\Lambda$, and cutting off and straightening; such a process would plausibly produce a Lagrangian satisfying the wrapping hypothesis. We contemplate this process because such asymptotically conical exact holomorphic Lagrangians appear frequently in examples, but do not literally provide objects of $Fuk(X,\partial_{\infty}\Lambda)$.
In a different direction, it was observed in [40] that the equivalence of microsheaf and Fukaya categories [11] remains true after enlarging the Fukaya category to contain unobstructed compact nonexact Lagrangians, of course taking coefficients in the Novikov field. As there, this observation is profitably combined with the fact [41] that holomorphic Lagrangians are unobstructed in hyperkähler manifolds. We conclude:
Corollary B.4.
Retain the hypotheses of Corollary B.2. Assume $W$ is hyperkähler and let $L\subset W$ be a compact holomorphic Lagrangian. Suppose $L\cup\Delta_{\alpha}$ is $g$-complex for each $\alpha$. Then $L\in Fuk(W,\partial_{\infty}\Lambda)^{\heartsuit}=\mu sh_{\Lambda}(\Lambda)^{\heartsuit}$, all categories taken with coefficients over the Novikov field.
Corollary B.3 and Corollary B.4 can be stated more generally as defining fully faithful functors from appropriate abelian category of local systems on $L$ to the heart of the $t$-structure.